OpenStudy (bbb911):

This is Geometry: In parallelogram ABCD, diagonals AC and BD intersect at point E. If AE = 4x-3, and CE = 17 -x, find x.

OpenStudy (anonymous):

AE=CE 4x-3=17-x 5x=20 x=4

OpenStudy (bbb911):

thanks again